# Metastability in the Dissipative Quantum Rabi Model

### v1

@@ IV Lifetime analysis and finite-size scaling
The principal component analysis provides us an rough estimation of the lifetime of the metastable parity-breaking state. To quantify this lifetime more accurately, we analyze the spectral decomposition of the Liouvillian $\mathcal{L}$ defined in Eq. (12). Taking the Liouvillian as a super-operator acting on the density matrix, the spectral decomposition of $\mathcal{L}$ is defined as
@@
where $\rho_{i}$ is the matrix representation of $|\rho_{i}\rangle\rangle$, and the coefficient $c_{i}=\langle\langle\tilde{\rho}_{i}|\rho(0)\rangle\rangle\equiv{\rm Tr}[\tilde{\rho}_{i}\rho(0)]$. Here, the sum of $i$ is split into two groups by using the criteria ${\rm Re}[\lambda_{m}]/{\rm Re}[\lambda_{m+1}]\ll 1$. The subspace spanned by the first $m$ eigenstates constitutes a metastable manifold, characterized by the fact that the eigenstates within it possess long lifetimes [Macieszczak2016]. The dimension of the metastable manifold is related with the symmetry of the Liouvillian. In the dissipative QRM, we have the parity symmetry, therefore only a pair of metastable states are supported, which gives $m=1$. The lifetime of the metastable state is then uniquely determined by the Liouvillian gap $\Delta\equiv{\rm Re}[\lambda_{1}]$ as
@@
For the parameters chosen in Fig. 4(c), the numerical calculation yields $\Delta/\omega_{0}=0.0034$, corresponding to $\omega_{0}T_{m}\approx 294.1$. This result is in quantitative agreement with that obtained via principal component analysis and quantum trajectory simulations.
Figure 7: (a) Finite-size scaling of the Liouvillian gap with fixing $\lambda=1.4\sqrt{{\Omega\omega_{0}}/{2}}$ and (b) the dependence of the Liouvillian gap $\Delta_{\infty}$ on the spin relaxation rate $\gamma$ in the thermodynamic limit. For (a), the numerical fitting $a+b\frac{\omega_{0}}{\Omega}+c\frac{\omega^{2}_{0}}{\Omega^{2}}$ yields coefficients $a=0.0034,b=0.068,c=1.54$ for $\gamma/\omega_{0}=0.05$ and $a=0.0,b=0.081,c=0.64$ for $\gamma/\omega_{0}=0$. For (b), the numerical fitting $a+b\frac{\gamma}{\omega_{0}}+c\frac{\gamma^{2}}{\omega_{0}^{2}}$ yields coefficients $a=0.0,b=0.066,c=0.055$. Other parameters are as same as those in Fig. 2.
The above numerical results are obtained for finite $\Omega/\omega_{0}$. It is necessary to examine the metastability of the parity-breaking solutions in the thermodynamic limit, defined by $\Omega/\omega_{0}\to\infty$ [Emary2003, Emary2003A]. In particular, we need to examine whether the Liouvillian gap closes in this limit. According to Liouvillian spectral theory, the gap closes in the symmetry-breaking phase in the thermodynamic limit [Minganti2018], since a symmetry-breaking steady state implies that all its symmetry-transformed counterparts are also steady states. Thus, the Liouvillian contains at least two eigenstates with zero eigenvalues. Conversely, if the Liouvillian gap remains open in the thermodynamic limit, the symmetry-breaking phase can only be metastable. A criterion of the SMP is then obtained as
@@
To assess the Liouvillian gap $\Delta_{\infty}$, we perform finite-size scaling with fixed $\lambda^{2}/\Omega\omega_{0}$. The results, presented in Fig. 7(a), show that the Liouvillian gap $\Delta$ decreases very slowly with increasing $\Omega/\omega_{0}$. A polynomial fit $\Delta=a+b\frac{\omega_{0}}{\Omega}+c\frac{\omega^{2}_{0}}{\Omega^{2}}$ yields that the gap approaches a finite value $\Delta_{\infty}\approx 0.0034$ as $\Omega/\omega_{0}\to\infty$, confirming the metastability of the parity-breaking states in the thermodynamic limit. For comparison, we also analyze the scaling of the gap at $\gamma=0$, which exhibits a power-law closing as $\Omega/\omega_{0}$ increases. This behavior indicates that the SP becomes stable in the absence of spin relaxation. However, the power-law closing, in contrast to the exponential closing of typical equilibrium phase transitions, suggests that the SP at $\gamma=0$ should still be regarded as metastable in a broader sense. A more precise characterization of the distinction between power-law and exponential gap closing is beyond the scope of this work. In Fig. 7(b), we show the dependence of the Liouvillian gap $\Delta_{\infty}$ on the spin relaxation rate $\gamma$, which reveals that $\Delta_{\infty}$ vanishes linearly with decreasing $\gamma$. This indicates that the SP is stable (metastable in a broader sense) only in the $\gamma=0$ limit. Even a small spin relaxation would render the SP to the SMP, which explains the phase diagram in Fig. 1(b).
@@ V Discussion
Metastability is a common feature in dissipative phase transitions [Haken1986, Letscher2017, Rota2018, Gerry2024]. For instance, in the lasing phase transition, the system transits from the normal phase to the lasing phase characterized by spontaneous breaking of the global U(1) symmetry above the lasing threshold. However, atomic pumping—which plays a role analogous to spin relaxation—continuously quench the system and introduces perturbations that prevent the laser maintaining a definite phase. This results in phase diffusion that gradually restores the global U(1) symmetry, rendering the lasing phase metastable.
@@
The stability analysis developed for the dissipative QRM can be generalized to other open quantum systems with dissipative phase transition. From this paper, we show that the linear stability analysis only characterizes the stability of the symmetry-breaking state under small perturbations. It is essential to assess whether quantum fluctuations are negligible, for example, by examining the convergence of the cumulant expansion. Another approach to evaluate the stability of the symmetry-breaking phase is to investigate the closing of the Liouvillian gap in the thermodynamic limit. However, for most many-body systems, calculating the Liouvillian gap is challenging, as in the case of lasing models. For such systems, the semiclassical Fokker–Planck equation can be used to estimate the phase diffusion rate, which is directly related to the Liouvillian gap.
@@
Through the study of the dissipative QRM, we can draw a comparison between phase transitions in equilibrium and open quantum systems. In equilibrium systems, phase transitions occur when the equilibrium state changes from a symmetry-preserving state to a symmetry-breaking state. This symmetry-breaking state typically corresponds to the global minimum of the free energy or the ground state energy. Although Goldstone modes exist as zero-energy excitations when a continuous symmetry is broken, the Mermin-Wagner theorem guarantees that such excitations can be neglected in high-dimensional systems [Mermin1966]. Therefore, the symmetry-breaking state in equilibrium systems is generally stable. In contrast, dissipative phase transition in open quantum systems exhibit fundamentally different behaviors. First, the steady state corresponds to local minima of an effective potential, which can give rise to multistable phases. Second, relaxation or driven processes—unavoidable in open systems—may act as random and strong quenches to the system. These processes can drive the system from one local minimum to another, making the resulting steady state being determined by a dynamical balance between different components in the steady state and leading to the metastability.
@@
Finally, a variety of experimental platforms can be used to verify the SMP of the dissipative QRM under a weak spin relaxation, such as trapped ions [Puebla2017, Lv2018, Cai2021], nuclear magnetic resonance  [Chen2021, Wu2024], and superconducting circuits [Braumueller2017, Langford2017]. Notably, the dissipative phase transition at $\gamma=0$ has been demonstrated in recent experiments [Wu2024]. We expect that the metastability of superradiant phase can be experimentally tested on these platforms.
@@ VI Conclusion
In this paper, we investigate the metastability of the dissipative phase transition of the QRM in the presence of a weak spin relaxation. We reveal that the symmetry-breaking states in this model are inherently metastable. Spin relaxation induces random spin flips, which further act as strong perturbations to the system, facilitating transitions between different symmetry-breaking states. These transitions ultimately restore the system symmetry and give rise to the metastability of the symmetry-breaking phase.
To elucidate the metastability of the dissipative QRM, we perform both mean-field and cumulant expansion analyses, as well as exact numerical simulations for finite $\Omega/\omega_{0}$. This reveal a one-to-one correspondence between the steady-state distribution and the dynamic evolution of the metastable state. Based on this correspondence, we identify the coexistence of different components in the SMP and provides a rough estimate of their relative lifetimes from the principal component analysis. To quantify the lifetime more precisely and explore the fate of the superradiant metastable phase in the thermodynamic limit, we analyze the Liouvillian spectrum. The results show that the lifetime is uniquely governed by the Liouvillian gap, and is also consistent with the principal component analysis. Finite-size scaling further confirms that the superradiant phase is stable only in the absence spin relaxation but becomes metastable once a small spin relaxation is present.
Metastable symmetry-breaking phases are a common feature of open quantum systems and underscore a fundamental distinction between phase transitions in equilibrium and non-equilibrium systems. Our findings offer new insights into the metastability of dissipative phase transitions in open quantum systems.
@@ Appendix A Cumulant expansion method
From the master equation in Eq. (2), one can derive the equation of motion for any operator $\hat{O}=\hat{o}^{n_{1}}_{i}\hat{o}^{n_{2}}_{j}\cdots$ (i.e., correlations) as
@@
with $\mathcal{D}^{\prime}_{\hat{a}}[\hat{O}]\equiv 2\kappa\hat{a}^{\dagger}\hat{O}\hat{a}-\kappa\hat{a}^{\dagger}\hat{a}\hat{O}-\kappa\hat{O}\hat{a}^{\dagger}\hat{a}$ and $\mathcal{D}^{\prime}_{\hat{\sigma}_{-}}[\hat{O}]\equiv 2\gamma\hat{\sigma}_{+}\hat{O}\hat{\sigma}_{-}-\gamma\hat{\sigma}_{+}\hat{\sigma}_{-}\hat{O}-\gamma\hat{O}\hat{\sigma}_{+}\hat{\sigma}_{-}$. The order of the correlations is defined as $n\equiv n_{1}+n_{2}+\cdots$, and the operators $\hat{o}_{i}$ can be $\hat{x}$, $\hat{p}$, or $\hat{\sigma}_{x/y/z}$ in the dissipative QRM. A specific example is
@@
It can be observed that the equations for second-order correlations couple to third-order correlations, giving rise to a hierarchy structure in which low-order correlations are coupled to higher-order ones. This hierarchy leads to an infinite set of coupled equations, thereby precluding an exact solution of the correlations.
Here, we employ the cumulant expansion method to truncate the hierarchy equations and close these correlation equations [Barquilla2020, Leymann2014, plankensteiner2022]. The cumulant $\langle\hat{o}^{n_{1}}_{i_{1}}\hat{o}^{n_{2}}_{i_{2}}\cdots\rangle_{c}$ is defined from the generating function $\langle e^{\sum_{l}\eta_{l}\hat{o}_{l}}\rangle$ as
@@
where $n=n_{1}+n_{2}+\cdots$ denotes the order of the cumulant [Kubo1962]. Through the definition of cumulants, one can establish relations between correlations and cumulants. For low-order correlations, we have
@@
One can observe that the $n$-th order correlations depend only on cumulants up to $n$-th order.
Applying the cumulant expansion to both sides of Eq. (25), we obtain the equations of motion for cumulants. The left-hand side includes correlations (and hence cumulants) up to $n$-th order, while the right-hand side includes correlations up to $(n+1)$-th order due to the hierarchy structure. By truncating the hierarchy through neglecting these $(n+1)$-th order cumulants on the right-hand side, the equations of motion for all operators $\{\hat{O}(t)\}$ up to $n$-th order become closed in terms of cumulants, enabling us to solve them.
A major advantage of the cumulant expansion method over naive truncation of correlations lies in its systematic treatment of genuine coupled correlations [Kubo1962]. Cumulants quantify the intrinsic correlations between operators, which generally decrease as the order increases. By neglecting the $(n+1)$-th order cumulants, all correlations up to the $n$-th order are retained exactly, while all higher-order correlations are approximated in a self-consistent manner. For example, the second-order truncation yields $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle\approx\langle\hat{o}_{i}\hat{o}_{j}\rangle_{c}\langle\hat{o}_{k}\rangle_{c}+\langle\hat{o}_{i}\hat{o}_{k}\rangle_{c}\langle\hat{o}_{j}\rangle_{c}+\langle\hat{o}_{i}\rangle_{c}\langle\hat{o}_{j}\hat{o}_{k}\rangle_{c}+\langle\hat{o}_{i}\rangle_{c}$ $\langle\hat{o}_{j}\rangle_{c}\langle\hat{o}_{k}\rangle_{c}$. This truncation provides a good approximation to $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle$ when $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle_{c}$ is negligible. Thus, the cumulant expansion method provides a controlled and physically motivated approximation that improves both the accuracy and stability of the resulting equations of motion.

### v2

@@ IV Lifetime analysis and finite-size scaling
The principal component analysis provides us a rough estimation of the lifetime of the metastable parity-breaking state. To quantify this lifetime more accurately, we analyze the spectral decomposition of the Liouvillian $\mathcal{L}$ defined in Eq. (12). Taking the Liouvillian as a super-operator acting on the density matrix, the spectral decomposition of $\mathcal{L}$ is defined as
@@
where $\rho_{i}$ is the matrix representation of $|\rho_{i}\rangle\rangle$ and the coefficient $c_{i}=\langle\langle\tilde{\rho}_{i}|\rho(0)\rangle\rangle\equiv{\rm Tr}[\tilde{\rho}_{i}\rho(0)]$. Here, the sum of $i$ is split into two groups by using the criterion ${\rm Re}[\lambda_{m}]/{\rm Re}[\lambda_{m+1}]\ll 1$. The subspace spanned by the first $m$ eigenstates constitutes a metastable manifold, characterized by the fact that the eigenstates within it possess long lifetimes [45]. The dimension of the metastable manifold is related to the symmetry of the Liouvillian. In the dissipative QRM, we the parity symmetry; therefore, only a pair of metastable states are supported, which gives $m=1$. The lifetime of the metastable state is then uniquely determined by the Liouvillian gap $\Delta\equiv{\rm Re}[\lambda_{1}]$ as
@@
For the parameters chosen in Fig. 4(c), the numerical calculation yields $\Delta/\omega_{0}=0.0035$, corresponding to $\omega_{0}T_{m}\approx 294.1$. This result is in quantitative agreement with that obtained via principal component analysis and quantum trajectory simulations.
Figure 7: (a) Finite-size scaling of the Liouvillian gap with $\lambda=1.4\sqrt{{\Omega\omega_{0}}/{2}}$ and (b) dependence of the Liouvillian gap $\Delta_{\infty}$ on the spin relaxation rate $\gamma$ in the thermodynamic limit. For (a), the numerical fitting $a+b\frac{\omega_{0}}{\Omega}+c\frac{\omega^{2}_{0}}{\Omega^{2}}$ yields coefficients $a=0.0034,b=0.068,c=1.54$ for $\gamma/\omega_{0}=0.05$ and $a=0.0,b=0.081,c=0.64$ for $\gamma/\omega_{0}=0$. For (b), the numerical fitting $a+b\frac{\gamma}{\omega_{0}}+c\frac{\gamma^{2}}{\omega_{0}^{2}}$ yields coefficients $a=0.0,b=0.066,c=0.055$. The other parameters are the same as those in Fig. 2.
The above numerical results are obtained for finite $\Omega/\omega_{0}$. It is necessary to examine the metastability of the parity-breaking solutions in the thermodynamic limit, defined by $\Omega/\omega_{0}\to\infty$ [17, 18]. In particular, we need to examine whether the Liouvillian gap closes in this limit. According to Liouvillian spectral theory, the gap closes in the symmetry-breaking phase in the thermodynamic limit [51], since a symmetry-breaking steady state implies that all its symmetry-transformed counterparts are also steady states. Thus, the Liouvillian contains at least two eigenstates with zero eigenvalues. Conversely, if the Liouvillian gap remains open in the thermodynamic limit, the symmetry-breaking phase can only be metastable. A criterion of the SMP is then obtained as
@@
To assess the Liouvillian gap $\Delta_{\infty}$, we perform finite-size scaling with fixed $\lambda^{2}/\Omega\omega_{0}$. The results, presented in Fig. 7(a), show that the Liouvillian gap $\Delta$ decreases very slowly with increasing $\Omega/\omega_{0}$. A polynomial fit $\Delta=a+b\frac{\omega_{0}}{\Omega}+c\frac{\omega^{2}_{0}}{\Omega^{2}}$ yields that the gap approaches a finite value $\Delta_{\infty}\approx 0.0034$ as $\Omega/\omega_{0}\to\infty$, confirming the metastability of the parity-breaking states in the thermodynamic limit. For comparison, we also analyze the scaling of the gap at $\gamma=0$, which exhibits a power-law closing as $\Omega/\omega_{0}$ increases. This behavior indicates that the SP becomes stable in the absence of spin relaxation. However, the power-law closing, in contrast to the exponential closing of typical equilibrium phase transitions, suggests that the SP at $\gamma=0$ should still be regarded as metastable in a broader sense. A more precise characterization of the distinction between power-law and exponential gap closing is beyond the scope of this work. In Fig. 7(b) we show the dependence of the Liouvillian gap $\Delta_{\infty}$ on the spin relaxation rate $\gamma$, which reveals that $\Delta_{\infty}$ vanishes linearly with decreasing $\gamma$. This indicates that the SP is stable (metastable in a broader sense) only in the $\gamma=0$ limit. Even a small spin relaxation would render the SP a SMP, which explains the phase diagram in Fig. 1(b).
@@ V Discussion
Metastability is a common feature in dissipative phase transitions [25, 37, 60, 24]. For instance, in the lasing phase transition, the system transits from the normal phase to the lasing phase characterized by spontaneous breaking of the global U(1) symmetry above the lasing threshold. However, atomic pumping, which plays a role analogous to spin relaxation, continuously quenches the system and introduces perturbations that prevent the laser maintaining a definite phase. This results in phase diffusion that gradually restores the global U(1) symmetry, rendering the lasing phase metastable.
@@
The stability analysis developed for the dissipative QRM can be generalized to other open quantum systems with dissipative phase transitions. In this paper we showed that the linear stability analysis only characterizes the stability of the symmetry-breaking state under small perturbations. It is essential to assess whether quantum fluctuations are negligible, for example, by examining the convergence of the cumulant expansion. Another approach to evaluate the stability of the symmetry-breaking phase is to investigate the closing of the Liouvillian gap in the thermodynamic limit. However, for most many-body systems, calculating the Liouvillian gap is challenging, as in the case of lasing models. For such systems, the semiclassical Fokker-Planck equation can be used to estimate the phase diffusion rate, which is directly related to the Liouvillian gap.
@@
Through the study of the dissipative QRM, we can draw a comparison between phase transitions in equilibrium and open quantum systems. In equilibrium systems, phase transitions occur when the equilibrium state changes from a symmetry-preserving state to a symmetry-breaking state. This symmetry-breaking state typically corresponds to the global minimum of the free energy or the ground state energy. Although Goldstone modes exist as zero-energy excitations when a continuous symmetry is broken, the Mermin-Wagner theorem guarantees that such excitations can be neglected in high-dimensional systems [49]. Therefore, the symmetry-breaking state in equilibrium systems is generally stable. In contrast, dissipative phase transition in open quantum systems exhibit fundamentally different behaviors. First, the steady state corresponds to local minima of an effective potential, which can give rise to multistable phases. Second, relaxation or driven processes, unavoidable in open systems, may act as random and strong quenches to the system. These processes can drive the system from one local minimum to another, making the resulting steady state determined by a dynamical balance between different components in the steady state and leading to the metastability.
@@
Finally, a variety of experimental platforms can be used to verify the SMP of the dissipative QRM under a weak spin relaxation, such as trapped ions [56, 43, 11], nuclear magnetic resonance  [12, 67], and superconducting circuits [9, 33]. Notably, the dissipative phase transition at $\gamma=0$ has been demonstrated in recent experiments [67]. We expect that the metastability of the superradiant phase can be experimentally tested on these platforms.
@@ VI Conclusion
In this paper we investigated the metastability of the dissipative phase transition of the QRM in the presence of a weak spin relaxation. We revealed that the symmetry-breaking states in this model are inherently metastable. Spin relaxation induces random spin flips, which further act as strong perturbations to the system, facilitating transitions between different symmetry-breaking states. These transitions ultimately restore the system symmetry and give rise to the metastability of the symmetry-breaking phase.
To elucidate the metastability of the dissipative QRM, we performed both mean-field and cumulant expansion analyses, as well as exact numerical simulations for finite $\Omega/\omega_{0}$. These revealed a one-to-one correspondence between the steady-state distribution and the dynamic evolution of the metastable state. Based on this correspondence, we identified the coexistence of different components in the SMP and provide a rough estimate of their relative lifetimes from the principal component analysis. To quantify the lifetime more precisely and explore the fate of the superradiant metastable phase in the thermodynamic limit, we analyzed the Liouvillian spectrum. The results show that the lifetime is uniquely governed by the Liouvillian gap and is also consistent with the principal component analysis. Finite-size scaling further confirmed that the superradiant phase is stable only in the absence of spin relaxation but becomes metastable once a small spin relaxation is present.
Metastable symmetry-breaking phases are a common feature of open quantum systems and underscore a fundamental distinction between phase transitions in equilibrium and nonequilibrium systems. Our findings offer insights into the metastability of dissipative phase transitions in open quantum systems.
@@ Appendix A Cumulant expansion method
From the master equation (2), we can derive the equation of motion for any operator $\hat{O}=\hat{o}^{n_{1}}_{i}\hat{o}^{n_{2}}_{j}\cdots$ (i.e., correlations) as
@@
with $\mathcal{D}^{\prime}_{\hat{a}}[\hat{O}]\equiv 2\kappa\hat{a}^{\dagger}\hat{O}\hat{a}-\kappa\hat{a}^{\dagger}\hat{a}\hat{O}-\kappa\hat{O}\hat{a}^{\dagger}\hat{a}$ and $\mathcal{D}^{\prime}_{\hat{\sigma}_{-}}[\hat{O}]\equiv 2\gamma\hat{\sigma}_{+}\hat{O}\hat{\sigma}_{-}-\gamma\hat{\sigma}_{+}\hat{\sigma}_{-}\hat{O}-\gamma\hat{O}\hat{\sigma}_{+}\hat{\sigma}_{-}$. The order of the correlations is defined as $n\equiv n_{1}+n_{2}+\cdots$, and the operators $\hat{o}_{i}$ can be $\hat{x}$, $\hat{p}$, or $\hat{\sigma}_{x,y,z}$ in the dissipative QRM. A specific example is
@@
It can be observed that the equations for second-order correlations couple to third-order correlations, giving rise to a hierarchical structure in which low-order correlations are coupled to higher-order ones. This hierarchy leads to an infinite set of coupled equations, thereby precluding an exact solution of the correlations.
Here we employ the cumulant expansion method to truncate the hierarchy equations and close these correlation equations [62, 38, 54]. The cumulant $\langle\hat{o}^{n_{1}}_{i_{1}}\hat{o}^{n_{2}}_{i_{2}}\cdots\rangle_{c}$ is defined from the generating function $\langle e^{\sum_{l}\eta_{l}\hat{o}_{l}}\rangle$ as
@@
where $n=n_{1}+n_{2}+\cdots$ denotes the order of the cumulant [1]. Through the definition of cumulants, one can establish relations between correlations and cumulants. For low-order correlations, we have
@@
One can observe that the $n$th order correlations depend only on cumulants up to $n$th order.
Applying the cumulant expansion to both sides of Eq. (25), we obtain the equations of motion for cumulants. The left-hand side includes correlations (and hence cumulants) up to $n$th order, while the right-hand side includes correlations up to $(n+1)$th order due to the hierarchy structure. By truncating the hierarchical through neglecting these $(n+1)$th order cumulants on the right-hand side, the equations of motion for all operators $\{\hat{O}(t)\}$ up to $n$th order become closed in terms of cumulants, enabling us to solve them.
A major advantage of the cumulant expansion method over naive truncation of correlations lies in its systematic treatment of genuine coupled correlations [1]. Cumulants quantify the intrinsic correlations between operators, which generally decrease as the order increases. By neglecting the $(n+1)$th order cumulants, all correlations up to the $n$th order are retained exactly, while all higher-order correlations are approximated in a self-consistent manner. For example, the second-order truncation yields $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle\approx\langle\hat{o}_{i}\hat{o}_{j}\rangle_{c}\langle\hat{o}_{k}\rangle_{c}+\langle\hat{o}_{i}\hat{o}_{k}\rangle_{c}\langle\hat{o}_{j}\rangle_{c}+\langle\hat{o}_{i}\rangle_{c}\langle\hat{o}_{j}\hat{o}_{k}\rangle_{c}+\langle\hat{o}_{i}\rangle_{c}$ $\langle\hat{o}_{j}\rangle_{c}\langle\hat{o}_{k}\rangle_{c}$. This truncation provides a good approximation to $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle$ when $\langle\hat{o}_{i}\hat{o}_{j}\hat{o}_{k}\rangle_{c}$ is negligible. Thus, the cumulant expansion method provides a controlled and physically motivated approximation that improves both the accuracy and stability of the resulting equations of motion.